OpenStudy (anonymous):

legit answer or am i way off base: Show that integral from 1 to infinity: (e^-x) / x dx converges. substitution gets you 0 / infinity L'hospital get you -e^-x/1 or converges to zero???

OpenStudy (anonymous):

\[\int\limits_{1}^{\infty}((e^-x)/x) dx\]

OpenStudy (amistre64):

Lhopital is for derivatives ...

OpenStudy (amistre64):

does the function -1/x^2 converge?

OpenStudy (anonymous):

yes by p test 2>1 and p>1 converges

OpenStudy (amistre64):

so lets integrate it; and we get the function 1/x does this converge?

OpenStudy (amistre64):

the convergence of a derivative is not a gaurentee that the integral will converge ...

OpenStudy (anonymous):

now p = 1 and by p test p=1 diverges...sorry slow wrote p test down wrong in class and was doing it backwards for three days

OpenStudy (anonymous):

so looks like i should integrate and see what that gets me...

OpenStudy (amistre64):

if we integrate this by parts; it might produce a series .....

OpenStudy (anonymous):

this will be slow if you want to help elsewhere...

OpenStudy (amistre64):

lol, its 130am .... im not really wanting to help anywhere :)

OpenStudy (anonymous):

so that looks like (x)(-e^(-x)) - e^(-x)

OpenStudy (amistre64):

\[\frac{1}{x}e^{-x}=\frac{1}{x}\sum_{n=0}^{inf}\frac{(-1)^{n+1}}{n!}x^n\] \[\frac{1}{x}e^{-x}=\sum_{n=0}^{inf}\frac{(-1)^{n+1}}{n!}x^{n-1}\]

OpenStudy (amistre64):

hmm, -1^n that is ... i seem to have generated a spurious negative :)

OpenStudy (anonymous):

if I got that then i would be thinking ratio test? and spurious sign changes is my signature move

OpenStudy (amistre64):

this is still not integrated; i know when we derive a summation we add to n each time; so i assume when we integrate we get a n-1 to the summantion since n needs to start at 0; lets pull out the first iteration before we integrate

OpenStudy (anonymous):

maybe that is what my key suggests...it just integrates the e^-x and the 1/x disappears in an unexplained poof of smoke

OpenStudy (amistre64):

\[\int \frac{1}{x}+\sum_{n=1}^{inf}\frac{(-1)^{n}}{n!}x^{n-1}\ dx\]

OpenStudy (amistre64):

\[ln(x) + \sum_{n=0}^{inf}\frac{(-1)^n}{n!(n)}x^n+C\]

OpenStudy (amistre64):

uh oh, we got a zero in the denominator with that one :)

OpenStudy (anonymous):

it will be ok if we start with 1

OpenStudy (amistre64):

yeah, but that ln(x) on the outside doesnt converge ...

OpenStudy (amistre64):

so i have a feeling this thing diverges. we could chk it with the wolf

OpenStudy (anonymous):

i think they disappeared the 1/x with the idea that if integrate e^-x and converge the multiply 1/x would just make it smaller?

OpenStudy (amistre64):

its possible i could have tried to modify it in another manner; i ahvent done enough of these to be too confident

OpenStudy (anonymous):

All this converge diverge stuff seems like cheating ....ignore that part just approximate...im just not sure when to do it

OpenStudy (anonymous):

well at least in know not to check the limits with l'hospital....seemed like such a good gambit too

OpenStudy (amistre64):

:) the summation we got converges by a ratio test .... if we turn the ln(x) into a summation (power series) and add the 2 together I wonder if it pans out

OpenStudy (anonymous):

im saving the series expansion for tomorrow...(part three of the final) I'll work the ratio test on that sum/series for practice thanks for the helps

OpenStudy (anonymous):

\[ \int_1^c \frac {e^{-x}}x dx \\ u=\frac {1}x, \, dv =e^{-x} dx\\ du =-\frac 1 {x^2},\, v =- e^{-x}\\ \int_1^c \frac {e^{-x}}x dx = \left [- \frac {e^{-x}}x \right]_1^c - \int_1^c \frac {e^{-x}}{x^2}dx=\\ - \frac {e^{-c}}c+ \frac {e^{-1}}1 -\int_1^c \frac {e^{-x}}{x^2}dx \]

OpenStudy (anonymous):

Notice that \[ \int_1^c \frac {e^{-x}} {x^2} dx \le \int_1^c \frac {dx} {x^2} = 1 - \frac 1 c\\ 0\le \int_1^\infty \frac {e^{-x}} {x^2} dx \le 1 \] Hence \[ \int_1^\infty \frac {e^{-x}} {x^2} dx \] is convergent. By this and my previous post the original integral is convergent.

OpenStudy (anonymous):

so that uses squeeze to prove convergence?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

If c goes to infinity \[ \int_1^\infty \frac {e^{-x}}x dx= \\ -0+ \frac {e^{-1}}1 - k \] Where \[ K = \int_1^\infty \frac {e^{-x}}{x^2}dx \]

OpenStudy (anonymous):

K=k

OpenStudy (anonymous):

oK

OpenStudy (anonymous):

Did you understand it?

OpenStudy (anonymous):

i dont think we did anything like that in class - you are subtracting the other integral from the original?

OpenStudy (anonymous):

oh ok that is from the integration by parts...now that makes sense excellent thanks

OpenStudy (anonymous):

yw